OpenStudy (anonymous):

write the equation of the line is slope intercept form that is parallel to y=2x-4 and contains the point (-6,10)

OpenStudy (abmon98):

Parallel lines share the same gradient y=2x-4 y=mx+c m=2 10=-6*2+c c=22 y=2x+22